# Exploring knowledge graph-based neural-symbolic system from application perspective

### v2

@@ Abstract
The rapid advancement in artificial intelligence (AI), particularly through deep neural networks, has catalyzed significant progress in fields such as vision and text processing. Nonetheless, the pursuit of AI systems that exhibit human-like reasoning and interpretability continues to pose a substantial challenge. The Neural-Symbolic paradigm, which integrates the deep learning prowess of neural networks with the reasoning capabilities of symbolic systems, presents a promising pathway toward developing more transparent and comprehensible AI systems. Within this paradigm, the Knowledge Graph (KG) emerges as a crucial element, offering a structured and dynamic method for representing knowledge through interconnected entities and relationships, predominantly utilizing the triple (subject, predicate, object). This paper explores recent advancements in neural-symbolic integration based on KG, elucidating how KG underpins this integration across three key categories: enhancing the reasoning and interpretability of neural networks through the incorporation of symbolic knowledge (Symbol for Neural), refining the completeness and accuracy of symbolic systems via neural network methodologies (Neural for Symbol), and facilitating their combined application in Hybrid Neural-Symbolic Integration. It highlights current trends and proposes directions for future research in the domain of Neural-Symbolic AI.
@@ 1 Introduction
With the rapid advancement of deep learning, particularly in deep neural networks (DNNs) within Artificial Intelligence (AI), we have observed the emergence of groundbreaking methods. These innovations have obtained significant achievements in fields such as vision and text processing. For instance, models like EfficientNet[1], ResNeSt[2], and Vision Transformer[3] have demonstrated exceptional performance in tasks like image classification, target detection, and image segmentation. Similarly, the NLP domain has seen substantial strides with deep neural network-based pre-trained language models, such as GPT-4[4], Llama 2[5], and BERT[6], setting new benchmarks in text comprehension and generation.
@@
The field of model interpretability offers a vast array of research avenues, with some scientists recently try to explore the concept of integrating neural and symbolic systems to address the black-box problem[20, 21, 22, 23, 24, 25, 26]. Yoshua Bengio, an ACM Turing Award laureate, highlighted in his 2019 NeurIPS presentation the necessity for deep learning to evolve from System 1 to System 2 thinking[27], where System 1 refers to the intuitive, fast, and unconscious cognitive processes that current deep learning[28] technologies excel in, and System 2 represents the logical, deliberate, and conscious cognitive processes, a hallmark of Symbolic artificial intelligence in the expert system stage[29, 30, 31], which employs explicit symbols and rules to emulate human logical reasoning. This concept of transition underpins the concept of neural-symbolic systems, aiming to marry the pattern recognition prowess of deep learning models with the structured knowledge representation and logical reasoning capabilities of symbolic logic systems, thereby offering efficient abilities in learning and generalization, and clear logic. (As shown in Table 1 shows the strengths and drawbacks of the neural and symbolic systems)
Also, knowledge graph (KG)[32], as an important member of symbolic logic, plays a crucial role in neural-symbolic integration. They are built on triples (subject, predicate, object) and form a graph structure that encapsulates real-world entities, concepts, and their interrelations. In neural symbolic systems, the KG not only serves as a repository of information but also acts as a bridge connecting symbolic logic and neural networks. It enriches the contextual information of the neural network and improves the decision-making and interpretability of the model, for example, in NLP tasks to help the network understand the relationship between words and entities. Meanwhile, symbolic reasoning relies on the logical rules and facts provided by the KG, which play an important role in both model training and reasoning. This idea of combining KG and neural symbolic integration enables us to build smarter, more reliable, and transparent AI systems.
@@ 2.2 Symbol for neural
In this context, "Symbol for Neural"[60, 61, 62, 63, 64, 65, 66, 67, 68, 69, 70, 71, 20, 72, 73, 74, 75, 76, 77, 78, 79, 80, 51, 52, 81, 55, 54, 82, 83, 79, 78, 84, 85, 83, 86, 78, 87, 84, 85], also known as "Reasoning for Learning", leverages symbolic systems like KG to furnish a priori knowledge and a logical framework, thereby guiding and shaping neural networks’ learning processes. Symbolic knowledge encoded in KG provides a rich source of structured information that enables neural networks to enhance their interpretability and decision-making capabilities. Moreover, KG serves not merely as passive repositories but as active participants, infusing neural networks with domain-specific rules and facts to bolster their learning efficiency. For instance, in developing algorithms for a recommendation system on an online education platform, a KG can categorize courses by content, difficulty, and progression, directing the neural network to tailor learning paths for users, thus optimizing learning outcomes. Figure 3 illustrates the architecture and rationale of this integrated approach.
@@ 2.3 Hybrid neural-symbolic integration
"Hybrid neural-symbolic integration"[88, 89, 90, 91, 92, 93, 94, 95, 96, 97, 98, 99, 100, 101, 102, 103, 104] shows a more dynamic way of interaction. In this approach, neural networks and symbolic reasoning complement, but are not subordinate to, each other, and work together to contribute to the comprehension and inference capabilities of the AI system. In this system, neural networks first process input data (e.g., images, text, etc.), extracting features and converting them into intermediate representations. These representations are then passed to a symbolic system, which utilizes this data for logical reasoning and possibly combines it with existing KG to make decisions or generate new knowledge. The results of the reasoning are not only used for direct decision-making output but are also fed back to the neural network to guide its further learning and parameter tuning, thus optimizing the performance of the overall system. Through this iterative feedback mechanism, the hybrid system can continuously optimize itself, and its components, the neural network and the symbolic system, can learn and adapt from each other’s processing results. This collaborative process grants the hybrid system resilience and adaptability, allowing it to efficiently manage complex tasks with accuracy and interpretability. Figure 4 illustrates the architecture of this integrated approach.
@@ 3.1.1 KG-based recommender systems
In the field of exploring KG-enhanced recommender systems, research has been divided into two main directions: path-based approaches[106, 107, 108], embedding-based[109, 81, 110] approaches and the two propagation-based approaches[52, 51] that we will introduce below.

### v3

@@ Abstract
Advancements in Artificial Intelligence (AI) and deep neural networks have driven significant progress in vision and text processing. However, achieving human-like reasoning and interpretability in AI systems remains a substantial challenge. The Neural-Symbolic paradigm, which integrates neural networks with symbolic systems, presents a promising pathway toward more interpretable AI. Within this paradigm, Knowledge Graphs (KG) are crucial, offering a structured and dynamic method for representing knowledge through interconnected entities and relationships, typically as triples (subject, predicate, object). This paper explores recent advancements in neural-symbolic integration based on KG, examining how it supports integration in three categories: enhancing the reasoning and interpretability of neural networks with symbolic knowledge (Symbol for Neural), refining the completeness and accuracy of symbolic systems via neural network methodologies (Neural for Symbol), and facilitating their combined application in Hybrid Neural-Symbolic Integration. It highlights current trends and proposes future research directions in Neural-Symbolic AI.
@@ 1 Introduction
With the rapid advancement of deep learning, particularly in deep neural networks (DNNs) within Artificial Intelligence (AI), we have observed the emergence of groundbreaking methods. These innovations have obtained significant achievements in fields such as vision and text processing. For instance, models like EfficientNet[1], ResNet[2], and Vision Transformer[3] have demonstrated exceptional performance in tasks like image classification, target detection, and image segmentation. Similarly, the NLP domain has seen substantial strides with deep neural network-based pre-trained language models, such as GPT-4[4], Llama 2[5], and BERT[6], setting new benchmarks in text comprehension and generation.
@@
The field of model interpretability offers a vast array of research avenues, with some scientists recently try to explore the concept of integrating neural and symbolic systems to address the black-box problem[20, 21, 22, 23, 24, 25, 26]. Yoshua Bengio, an ACM Turing Award laureate, highlighted in his 2019 NeurIPS presentation the necessity for deep learning to evolve from System 1 to System 2 thinking[27]. System 1 refers to the intuitive, fast, and unconscious cognitive processes that current deep learning[28] technologies excel in. In contrast, System 2 represents the logical, deliberate, and conscious cognitive processes, a hallmark of Symbolic artificial intelligence in the expert system stage[29, 30, 31]. This stage employs explicit symbols and rules to emulate human logical reasoning. This concept of transition underpins the concept of neural-symbolic systems, aiming to marry the pattern recognition prowess of deep learning models with the structured knowledge representation and logical reasoning capabilities of symbolic logic systems, thereby offering efficient abilities in learning and generalization, and clear logic. (As shown in Table 1 shows the strengths and drawbacks of the neural and symbolic systems)
Knowledge graph (KG)[32], as an important member of symbolic logic, plays a crucial role in neural-symbolic integration. They are built on triples (subject, predicate, object) and form a graph structure that encapsulates real-world entities, concepts, and their interrelations. In neural symbolic systems, the KG not only serves as a repository of information but also acts as a bridge connecting symbolic logic and neural networks. It enriches the contextual information of the neural network and improves the decision-making and interpretability of the model, for example, in NLP tasks to help the network understand the relationship between words and entities. Meanwhile, symbolic reasoning relies on the logical rules and facts provided by the KG, which play an important role in both model training and reasoning. This idea of combining KG and neural symbolic integration enables us to build smarter, more reliable, and transparent AI systems.
@@ 2.2 Symbol for neural
In this context, "Symbol for Neural"[60, 61, 62, 63, 64, 65, 66, 67, 68, 69, 70, 71, 20, 72, 73, 74, 75, 76, 77, 78, 79, 80, 51, 52, 81, 55, 54, 82, 83, 79, 78, 84, 85, 83, 86, 78, 87, 84, 85], also known as "Reasoning for Learning", leverages symbolic systems like KG to furnish a prior knowledge and a logical framework, thereby guiding and shaping neural networks’ learning processes. Symbolic knowledge encoded in KG provides a rich source of structured information that enables neural networks to enhance their interpretability and decision-making capabilities. Moreover, KG serves not merely as passive repositories but as active participants, infusing neural networks with domain-specific rules and facts to bolster their learning efficiency. For instance, in developing algorithms for a recommendation system on an online education platform, a KG can categorize courses by content, difficulty, and progression, directing the neural network to tailor learning paths for users, thus optimizing learning outcomes. Figure 3 illustrates the architecture and rationale of this integrated approach.
@@ 2.3 Hybrid neural-symbolic integration
"Hybrid neural-symbolic integration"[88, 89, 90, 91, 92, 93, 94, 95, 96, 97, 98, 99, 100, 101, 102, 103, 104] shows a more dynamic way of interaction. In this approach, neural networks and symbolic reasoning complement each other without being subordinate, working together to enhance the comprehension and inference capabilities of the AI system. In this system, neural networks first process input data (e.g., images, text, etc.), extracting features and converting them into intermediate representations. These representations are then passed to a symbolic system, which utilizes this data for logical reasoning and possibly combines it with existing KG to make decisions or generate new knowledge. The results of the reasoning are not only used for direct decision-making output but are also fed back to the neural network to guide its further learning and parameter tuning, thus optimizing the performance of the overall system. Through this iterative feedback mechanism, the hybrid system can continuously optimize itself, and its components, the neural network and the symbolic system, can learn and adapt from each other’s processing results. This collaborative process grants the hybrid system resilience and adaptability, allowing it to efficiently manage complex tasks with accuracy and interpretability. Figure 4 illustrates the architecture of this integrated approach.
@@ 3.1.1 KG-based recommender systems
In the field of exploring KG-enhanced recommender systems, research has been divided into three main directions: path-based approaches[106, 107, 108], embedding-based[109, 81, 110] approaches and propagation-based approaches[52, 51] that we will introduce below.